OpenStudy (anonymous):

Determine the solution set of (x + 2)2 = 36. {6, -8} {4, -8} {4, -6}

OpenStudy (anonymous):

First distribute the 2 and what do you get?

OpenStudy (anonymous):

how

OpenStudy (anonymous):

Take the 2 and distribute it to (x+2) and you get 2x+4. You multiply it

OpenStudy (anonymous):

After you end up with 2x+4=36. Subtract 4 on both sides, what's left?

OpenStudy (anonymous):

2-4

OpenStudy (anonymous):

is that wat you mean

OpenStudy (anonymous):

Don forget about the x's. when you minus 4 on both sides, the four on the left cancels out and then you subtract 36-4=?

OpenStudy (anonymous):

32

OpenStudy (anonymous):

Okay so the leftover equation is 2x=32. Divide by 2 on both sides to get x=

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Okay so you have x=16. Plug that into the original equation to get y

OpenStudy (anonymous):

explian please

OpenStudy (anonymous):

how do i get y

OpenStudy (anonymous):

Plug 16 into the original equation

OpenStudy (anonymous):

(16+2)2=36